# WRDScore: New Metric for Evaluation of Natural Language Generation Models

### v1

@@ Abstract.
software engineering, recommender systems, method name prediction
1. Introduction
The current rapid rise of deep-learning-based natural language generation models acutely necessitates the existence of smart metrics to evaluate these emerging models. Nowadays, the available metrics are unable to capture more subtle differences in semantics and syntax between generated and reference texts. That is, such simple and widely accepted metrics as ROUGE (Lin, 2004) can measure only the direct overlap between the tokens in reference and generated sequences. This naive approach does not consider possible paraphrasing, replacement of some tokens with synonyms, abbreviations, the order of tokens, etc. For instance, ROUGE-1 metric would not be able to detect the obvious similarity between words “copy” and “clone”, or “size” and “get count”.
In this paper, our goal is to create a novel metric for evaluation of the output of natural language generation (NLG) models, which is better tailored to the prediction problem of Java method names. Other embedding-based NLG metrics, such as BERTScore (Zhang et al., 2020), make unnecessary implicit assumptions. For instance, in the framework of BERTScore, one token from the reference sequence can be matched only to a single token in the predicted sequence, and vice versa. In this paper, we will show that this assumption can not only be safely relaxed, but leads to the considerable improvement even without using the advanced deep contextualized embeddings, such as BERT (Devlin et al., 2019).
@@ 1. Introduction
2. Related work
2.1. ROUGE
ROUGE metric (Lin, 2004), or Recall-Oriented Understudy for Gisting Evaluation, is the simplest and most widely used approach to compare two text sentences. It relies on the direct overlap between the two token sets of the input sentences. The most popular variation of ROUGE is the ROUGE-N metric, with N equal to 1 or 2. More precisely, ROUGE-N measures the number of matching n-grams the model-generated text and a human-produced reference sentence. ROUGE-N is a recall-oriented measure defined as follows:
$ROUGE-N=\frac{\sum_{S\in\{ReferenceSummaries\}}\sum_{gram_{n}\in S}Count_{match}(gram_{n})}{\sum_{S\in\{ReferenceSummaries\}}\sum_{gram_{n}\in S}Count(gram_{n})}$
Here $n$ stands for the length of the n-gram, $gram_{n}$, and $Count_{match}(gram_{n})$ is the maximum number of n-grams co-occurring in a candidate summary and a set of reference summaries. A closely related measure, BLEU, used in automatic evaluation of machine translation, is a precision-based measure. BLEU measures how well a candidate translation matches a set of reference translations by counting the percentage of n-grams in the candidate translation overlapping with the references.
2.2. BERTScore
BERTScore (Zhang et al., 2020) uses BERT (Devlin et al., 2019) contextualized embeddings to compute pairwise cosine similarity between each hypothesis token $i$ with each reference token $j$. Then, the precision and recall are calculated by maximum-based hard matching of tokens in the hypothesis sentence to the tokens of the reference sentence as follows:
@@ 2.2. BERTScore
2.3. MoverScore
MoverScore (Zhao et al., 2019) was inspired by using the Word Mover’s Distance (Kusner et al., 2015) (WMD) and takes the value equal to the cost of the optimal transportation plan between the reference and predicted sentences when they are viewed as probability distributions. MoverScore is able to use Word2Vec, BERT, or ELMo to define word embeddings. Also, MoverScore assigns weights for the defined word vectors via TF-IDF. Word embeddings can be n-grams.
It was observed that MoverScore (Zhao et al., 2019) only marginally outperforms BERTScore when measured on a variety of tasks, such as machine translation, image captioning, and dialog response generation. MoverScore outperforms both asymmetric HMD (“Hard Mover’s Distance”) factors (Zhao et al., 2019). However, if they are combined via harmonic mean, BERTScore is on par with MoverScore. We conjecture that BERT softens hard alignments of BERTScore as contextualized embeddings encode information about the whole sentence into each word vector.
@@ 3. Proposed Algorithm
This simple cooccurrence-based approach for building token embeddings is not the only choice. Some more advanced methods can be also be employed, such as Word2Vec (!!!cite!!!) or BERT. During our experiments, we observed that training Word2Vec over the training corpus with the minimal context window size, which is 2, results in embeddings that are able to capture synonym relations between tokens most precisely. This leads to a better overall performance of WRDScore.
@@
(1) $PR=\frac{1}{|p|}\sum_{i\in p}\max_{j\in r}\vec{i}^{T}\cdot\vec{j}$
(2) $RC=\frac{1}{|r|}\sum_{j\in r}\max_{i\in p}\vec{i}^{T}\cdot\vec{j}$
@@
To overcome the above limitation, we propose to use Word Rotator’s Distance (Yokoi et al., 2020). Now, suppose that the reference and predicted sequences are equipped with some set of weights $W=\{w_{1},\ldots,w_{l}\}$ and $\hat{W}=\{\hat{w}_{1},\ldots,\hat{w}_{m}\}$, respectively. The embedding norms of the corresponding tokens are usually taken as weights. This choice is justified since the norm of a token embedding is usually proportional to the token’s importance in the training corpus. These weights can also be augmented with the TF-IDF coefficients. Next, let $\{F_{ij}\}$, $i=1,\ldots,l$, $j=1,\ldots,m$, be the flow matrix which was the result of the solution of the following optimization problem:
(3) $\min_{F\in\mathbb{R}^{l\times m}}\sum_{i,j}F_{ij}\cdot d_{ij}$
@@
(4) $PR=\frac{1}{|p|}\sum_{i\in p}\frac{\sum_{j\in r}F_{ij}\cdot\cos(t_{i},\hat{t}_{j})}{\sum_{j\in r}F_{ij}}$
(5) $RC=\frac{1}{|r|}\sum_{i\in r}\frac{\sum_{j\in p}F_{ij}\cdot\cos(\hat{t}_{j},t_{i})}{\sum_{j\in p}F_{ij}}$
@@
4. Experiments
To demonstrate the efficiency of the proposed metric, we conducted some experiments on the following large dataset of Java code. By training the embeddings via the simple cooccurrence model with the largest context window size on Java method names from this dataset, we evaluate the output of an NLG model, called UniXcoder (Guo et al., 2022), and obtain the following results. Each line below represents the reference and predicted sequences as well as the new recall, precision, and F1 scores, for which the old ROUGE-1 score was zero:
  1. (1)
@@ 4. Experiments
Table 1. The observed mean deviation of the ground truth (human scores) from the scores obtained by the ROUGE-1 and WRDScore metrics
@@
5. Conclusion

### v3

@@ Abstract
Index Terms:
natural language generation, method name prediction, AI for code
I Introduction
The current rapid rise of deep-learning-based natural language generation models acutely necessitates the existence of smart metrics to evaluate these emerging models. Nowadays, the available metrics are unable to capture more subtle differences in semantics and syntax between generated and reference texts. That is, such simple and widely accepted metrics as ROUGE [1] can measure only the direct overlap between the tokens in reference and generated sequences. This naive approach does not consider possible paraphrasing, replacement of some tokens with synonyms, abbreviations, the order of tokens, etc. For instance, ROUGE-1 metric would not be able to detect the obvious similarity between words “copy” and “clone”, or “size” and “get count”.
In this paper, our goal is to create a novel metric for evaluation of the output of natural language generation (NLG) models, which is better tailored to the prediction problem of Java method names. Other embedding-based NLG metrics, such as BERTScore [4], make unnecessary implicit assumptions. For instance, in the framework of BERTScore, one token from the reference sequence can be matched only to a single token in the predicted sequence, and vice versa. In this paper, we will show that this assumption can not only be safely relaxed, but leads to the considerable improvement even without using the advanced deep contextualized embeddings, such as BERT [6].
@@ I Introduction
II Related work
II-A ROUGE
ROUGE metric [1], or Recall-Oriented Understudy for Gisting Evaluation, is the simplest and most widely used approach to compare two text sentences. It relies on the direct overlap between the two token sets of the input sentences. The most popular variation of ROUGE is the ROUGE-N metric, with N equal to 1 or 2. More precisely, ROUGE-N measures the number of matching n-grams in the model-generated text and the human-produced reference sentence. ROUGE-N is a recall-oriented measure defined as follows:
[…]
where $R$ referes to the set of reference summaries, $N$ stands for the length of the n-gram ($gram_{n}$), and $Count_{match}(gram_{n})$ is the maximum number of n-grams co-occurring in the candidate summary and the set of reference summaries. A closely related measure, BLEU, used in automatic evaluation of machine translation, is a precision-based measure. BLEU measures how well a candidate translation matches a set of reference translations by counting the percentage of n-grams in the candidate translation overlapping with the references.
II-B BERTScore
BERTScore [4] uses BERT [6] contextualized embeddings to compute pairwise cosine similarity between each hypothesis token $i$ with each reference token $j$. Then, the precision and recall are calculated by maximum-based hard matching of tokens in the hypothesis sentence to the tokens of the reference sentence as follows:
@@ II-B BERTScore
II-C MoverScore
MoverScore [5] was inspired by using the Word Mover’s Distance [7] (WMD) and takes the value equal to the cost of the optimal transportation plan between the reference and predicted sentences when they are viewed as probability distributions. MoverScore is able to use Word2Vec, BERT, or ELMo to define word embeddings. Also, MoverScore assigns weights for the defined word vectors via TF-IDF. Word embeddings can be n-grams.
It was observed that MoverScore [5] only marginally outperforms BERTScore when measured on a variety of tasks, such as machine translation, image captioning, and dialog response generation. MoverScore outperforms both asymmetric HMD (“Hard Mover’s Distance”) factors [5]. However, if they are combined via harmonic mean, BERTScore is on par with MoverScore. We conjecture that BERT softens hard alignments of BERTScore as contextualized embeddings encode information about the whole sentence into each word vector.
@@ III Proposed Algorithm
This simple cooccurrence-based approach for building token embeddings is not the only choice. Some more advanced methods can be also be employed, such as Word2Vec [8] or BERT. During our experiments, we observed that training Word2Vec over the training corpus with the minimal context window size, which is 2, results in embeddings that are able to capture synonym relations between tokens most precisely. This leads to a better overall performance of WRDScore.
@@
$PR=\frac{1}{|p|}\sum_{i\in p}\max_{j\in r}\vec{i}^{T}\cdot\vec{j}$ (1)
$RC=\frac{1}{|r|}\sum_{j\in r}\max_{i\in p}\vec{i}^{T}\cdot\vec{j}$ (2)
@@
To overcome the above limitation, we propose to use Word Rotator’s Distance [2]. Now, suppose that the reference and predicted sequences are equipped with some set of weights $W=\{w_{1},\ldots,w_{l}\}$ and $\hat{W}=\{\hat{w}_{1},\ldots,\hat{w}_{m}\}$, respectively. The embedding norms of the corresponding tokens are usually taken as weights. This choice is justified since the norm of a token embedding is usually proportional to the token’s importance in the training corpus. These weights can also be augmented with the TF-IDF coefficients. Next, let $\{F_{ij}\}$, $i=1,\ldots,l$, $j=1,\ldots,m$, be the flow matrix which was the result of the solution of the following optimization problem:
$\min_{F\in\mathbb{R}^{l\times m}}\sum_{i,j}F_{ij}\cdot d_{ij}$ (3)
@@
$PR=\frac{1}{|p|}\sum_{i\in p}\frac{\sum_{j\in r}F_{ij}\cdot\cos(t_{i},\hat{t}_{j})}{\sum_{j\in r}F_{ij}}$ (4)
$RC=\frac{1}{|r|}\sum_{i\in r}\frac{\sum_{j\in p}F_{ij}\cdot\cos(\hat{t}_{j},t_{i})}{\sum_{j\in p}F_{ij}}$ (5)
@@
IV Experiments
To demonstrate the efficiency of the proposed metric, we conducted some experiments on the following large dataset of Java code. By training the embeddings via the simple cooccurrence model with the largest context window size on Java method names from this dataset, we evaluate the output of an NLG model, called UniXcoder [3], and obtain the following results. Each line below represents the reference and predicted sequences as well as the new recall, precision, and F1 scores, for which the old ROUGE-1 score was zero:
  1. 1.
@@ IV Experiments
TABLE I: The observed mean deviation of the ground truth (human scores) from the scores obtained by the ROUGE-1 and WRDScore metrics
@@
V Reproducibility Package
The reproducibility package for this article is available at GitHub: https://github.com/rmusab/wrd-score.
VI Conclusion